OpenStudy (anonymous):

The square of a number decreased by 25 is -9. How do I find the x intercepts and the vertex of the graph

OpenStudy (anonymous):

x intercepts: set equation to zero and solve for x. vertex: put equation into standard quadratic form and use formula -b/2a, then plug value you get into function and get y coordinate

OpenStudy (anonymous):

Can yous how me?

OpenStudy (anonymous):

@alanli123

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

\[x ^{2}-25=-9\]

OpenStudy (anonymous):

Right then \[x ^{2} - 25 + 9 = 0\]

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

What happens next?

OpenStudy (anonymous):

Solve for x.

OpenStudy (anonymous):

\[x ^{2}=16\]

OpenStudy (anonymous):

and then you square them both so x = 4 right?

OpenStudy (anonymous):

x = postive AND minus 4

OpenStudy (anonymous):

ooh right cuz you squared it, what do we do next?

OpenStudy (anonymous):

Wait we kinda skipped something. We have to set the equation to 0, so: \[x ^{2}-16=0\] Then difference of perfect squares and factor it out. \[(x-4)(x+4)=0\] Solving the equation gives x equals plus or minus 4. (I

OpenStudy (anonymous):

We did this because this is the correct approach to solving the equation. The way before was a bit unorthodox.

OpenStudy (anonymous):

ohh

OpenStudy (anonymous):

So, that means the x intercepts is 4,0 and -4.0

OpenStudy (anonymous):

Now, for the vertex. So the quadratic equation we see so far is \[x ^{2}-16 = 0\] Since the coefficient in front of \[x^{2}\] is positive, that means the vertex is a minimum value one. Now, lets think about this with a graphical approach. x^2 - 16 = 0 is just x^2, except 16 units down, right? Well, the minimum value or vertex of x^2 is just 0, so that means if we subtract 16, we obtain the minimum value or vertex of x^2 - 16. So, the vertex of that is 0, -16

OpenStudy (anonymous):

Thanks a lot, know your tired (saw it in the chat) so thanks for helping me out

OpenStudy (anonymous):

haha np